# Quasar Pairs as Large–Scale Structure tracers

### v1

@@ 4 Quasar Pair Catalog
The distribution of distances in Mpc for all pairs is illustrated in Fig. (5). A total of 618 pairs were detected. In particular, the average distance between pairs is given by $d=(1.05\pm 0.51)\text{ Mpc}$, indicating that a significant portion of the distribution ($50.97\%$) lies beyond $d=1.05\text{Mpc}$. In short, the peak of the histogram suggests that the quasar pairs exhibit a preference for being situated within this specific projected distance. We also examine various bounds of $|\Delta V|$ to investigate how the number of pairs detected changes as this condition is relaxed. For condition $|\Delta V|\leq 600\text{km}/s$, the number of pairs detected decreases to 418, representing $0.29\%$ of the total. In contrast, when we consider $|\Delta V|\leq 1200\text{km}/s$, the number of pairs detected increases to 779, which represents $0.54\%$ of the entire subsample. Consequently, we will focus on the most conservative case moving forward, which corresponds to $|\Delta V|\leq 900\text{km}/s$. The distribution of quasar pairs within the RA-DEC plane is illustrated in Fig. (6) for the conservative scenario.
Figure 6: For the case where $|\Delta V|\leq 900\text{km}/s$, the scatter plot of the quasar pairs is presented.
It is now crucial to evaluate whether these quasar pairs exhibit certain similarities that classify them as distinct physical objects. We will start by calculating the bolometric luminosity of the sub-sample. The subsequent analysis will enable us to quantify the total energy emitted by these quasars Wu&Shen (2022), as well as the relationship between the luminosities of the pairs.
To compute the bolometric luminosity, we begin by considering the apparent magnitude in the I-band, denoted $m_{i}$. We can estimate the corresponding absolute magnitude $M_{i}$ using the redshift of the quasars, according to Planck18 fiducial cosmology. To do so, we utilize the distance modulus relation:
$M_{i}=m_{i}-5\log_{10}\left[D_{L}(z)\right]+15,$ (3)
where $D_{L}(z)$ represents the luminosity distance in megaparsecs units. Once we have $M_{i}$, we can convert it to bolometric luminosity in solar units. The conversion formula is expressed as follows:
$L_{bol}=10^{-0.4(M_{i}-M_{\odot,i})}L_{\odot},$ (4)
@@
where $M_{\odot,i}=4.08$ is the absolute magnitude of the Sun in the I-band, and $L_{\odot}$ is the solar luminosity. The latter approach allows us to derive a robust estimate of the bolometric luminosity for our subsample of quasars, enabling further analysis of their properties and potential physical similarities for the quasar pairs. In Fig. (7), the distribution of luminosities is displayed, showing that $L_{bol}=(5.08\pm 7.12)\times 10^{14}L_{\odot}$.
Figure 7: Histogram for the quasar distribution of luminosities
Our next step involves the examination of two distinct types of data sets. On the one hand, we will analyze the distribution of luminosity ratios, specifically $L_{1}/L_{2}$, corresponding to the quasar pairs, where $L_{1}$ represents the smaller value of the two luminosities, and $L_{2}$ denotes the larger one. We are able to investigate the intrinsic relationships between the luminosities of these objects, providing insight into their relative behaviors. On the other hand, we will construct 100 random sets of fake pairs derived from the isolated set  [^7] A comparable analysis was conducted on 250 random sets of synthetic pairs generated from the isolated set, revealing no significant deviations.. That methodology will enable us to establish a baseline for comparison, which is crucial to understanding the underlying characteristics of the quasar pairs and their luminosity distributions. By juxtaposing the observational data with these synthetic constructs, we aim to discern any significant patterns or anomalies that may emerge, thereby enriching our comprehension of these astrophysical entities.
Figure 8: Distribution of the luminosities ratios for true pairs and fake ones.
The physical criteria for the assembly of the artificial pairs stipulate that the angular separation between the objects must exceed $5.7^{\circ}$ (1 radian), and the maximum allowable redshift difference is restricted to $\delta z\leq 0.3$. In constructing these fake pairs, we ensured that there was no double counting, thus maintaining the integrity of our data set. We generated a total of 100 sets of fake pairs, each containing the same number of detected pairs, resulting in each set comprising 618 artificial pairs. We pay particular attention to the assembly of the new set by employing a random algorithm designed to avoid any selection bias.
Fig.(8) presents compelling evidence contrasting the characteristics of the true pairs against those of the fake pairs. In particular, the mean luminosity ratio $L_{1}/L_{2}$ demonstrates statistically significant differences between the two data sets. Furthermore, when we compute the fraction of real pairs that meet the condition $F_{\text{true pair}}(L_{1}/L_{2}<0.3)$, and compare it with the corresponding fraction for the artificial pairs $F_{\text{fake pair}}(L_{1}/L_{2}<0.3)$, we find that the ratio relation ${\cal{R}}=F_{\text{true pair}}/F_{\text{fake pair}}<1$ is true. The latter observation indicates that the two sets of quasar pairs exhibit fundamentally different properties, probably attributable to the different accretion mechanisms that influence the luminosity ratios $L_{1}/L_{2}$. Such differences underscore the complexity of these astrophysical systems and suggest that the evolution of these objects and the physical processes that govern their emissions are not uniform across the sample. It is important to note that our finding remains consistent across all 100 datasets of fake pairs, each composed of 618 pairs.
@@
Figure 9: Distribution of the logarithm of the sum of luminosities for true pairs and artificial pairs.
Another approach to investigate the validity of random-generated data sets compared to true pairs is to analyze the behavior of the logarithm of the sum of the luminosities. The rationale behind further inquiry lies in the possibility of encountering pairs where the luminosity ratio $L_{1}/L_{2}$ exhibits specific patterns; for example, both quasars may possess either low or high luminosities. Such a scenario raises the question of how to effectively discriminate between extreme cases. By examining the logarithmic sum of the luminosities, insight can be gained into the collective luminosity characteristics of each pair, allowing for the identification of trends that differentiate genuine pairs from synthetic counterparts. As a proof of concept, one specific comparison is presented; however, it is essential to emphasize that the approach holds true for all 100 dataset realizations. Fig. (9) clearly reveals that the logarithm of the sum of the luminosities does not adhere to the same distribution for true pairs and artificial pairs. This discrepancy further highlights the fundamental differences between the two datasets, suggesting that the underlying mechanisms driving the luminosity characteristics of quasars are inherently distinct. Differences may primarily stem from the co-evolution of quasars with their surrounding environments, influencing their accretion processes and, consequently, their luminosity profiles Porciani (2004). Such interactions with the environment can play a crucial role in shaping the properties of quasars, leading to observable variations in luminosity distributions Eftekharzadeh (2015).
@@
Figure 10: The distribution of luminosity ratios for true pairs and simulated pairs is illustrated.
We aim to further quantify the observed differences between true quasar pairs and their fake counterparts. The central premise is to evaluate whether a modified set of ”fake” pairs, now referred to as ”simulated” quasar pairs, can effectively approximate the real data. In this approach, we focus on the tercile of the fake sample associated with the less luminous quasar, where we increase its luminosity by $30\%$. Fig. (10) illustrates that this adjustment will produce a histogram of the luminosity ratio $L_{1}/L_{2}$ that more closely resembles the actual data (similarity effect). Indeed, we obtain a histogram that resembles the true data pairs; while not identical, it is quite close. For example, the true pairs have a ratio of $L_{1}/L_{2}=(0.472\pm 0.011)$, while the simulated pairs produce $L_{1}/L_{2}=(0.439\pm 0.011)$, resulting in a mean value difference of less than $7\%$. In a fundamental sense, this observation suggests that the true pairs exhibit closely aligned luminosity values. In contrast, fake pairs demonstrate a more pronounced disparity between the most luminous and the least luminous members. This distinction may imply that the true pairs share a common cosmic evolutionary origin, while the fake pairs do not.
@@
Figure 11: Top panel: Color-color map for the entire subsample of quasars. Bottom panel: Color-color map for the detected quasar pairs ($\text{QSO}_{1}-\text{QSO}_{2}$). Here, $\log(N)$ represents the logarithm of the number of quasar pairs that fall within each region of the color-color map. Notice that the maximum or minimum values may be outside the range of displayed data points in some cases.
@@
In the final stage of characterizing the quasar pairs, we perform a comparative analysis of the color-color map of these pairs against that of the entire selected sample. As illustrated in Fig. (11), the color-color map derived from the PSF magnitudes in the $g$, $i$, $r$ and $z$ bands for the complete subsample reveals a range of values given by $-6.04\leq(g-r)\leq 5.64$ and $-3.68\leq(r-z)\leq 7.15\leavevmode\nobreak$ Lyke (2020). However, when we focus on the detected pairs, these maximum and minimum values are significantly constrained, $-1.55\leq(g-r)\leq 2.21$ and $-1.17\leq(r-z)\leq 2.53$. It is important to note that $90\%$ of the data points associated with the quasar pairs fall within the interval $(r-z)\times(g-r)\in[-0.10,0.85]\times[-0.17,0.42]$. The latter results indicate older stellar populations or dust effects along with the possibility of bluer colors signifying ongoing star formation. This is consistent with the current understanding of quasar environments and their host galaxies Kauffmann &\&& Heckman (2009). For the $90\%$ of the detected quasar pairs, we obtain $(i-z)\times(r-i)\in[-0.06,0.69]\times[-0.13,0.30]$. Once again, the negative minimum values suggest the presence of younger, hotter stars, while the positive maximum values indicate the influence of older stars or dust obscuration. These interpretations are well supported within the framework of quasar studies Richards (2006). Different sequences of color-color maps are displayed in Fig. (12)
@@ 5 Cross-correlation Quasar-Pair
To ensure that the identified pairs are indeed physical composite objects located within the same dark matter halo and are likely experiencing similar accretion processes—rather than being merely coincidental nearby quasars—we will implement a series of verification tests. Among these, we will conduct an analysis of the cross-correlation of signals between the quasar sample and the corresponding subsample of pairs, as we stated before. We implement the same protocol as previously done. Using the corrfunc package Sinha (2020), we calculate $\xi(s,\mu)$, subsequently converting this to the number of pairs with the assistance of the convert-3d-counts-to-cf routine. That enables us to derive the Landy & Szalay (1993) estimator for the cross-correlation,
@@
where $D_{q}$ represents the set of quasars belonging to the subsample, and $D_{p}$ denotes the dataset of the quasar pairs, using the mid-vector between the two quasars to represent each pair. We implement a method to generate an associated random set that satisfies the specifications outlined in the corrfunc manual Sinha (2023). Specifically, we ensure that the size of the random set and the data set is consistent with the relationship $N(R_{p})=3N(D_{p})$. The latter approach improves statistical robustness and mitigates any potential biases in our correlation measurements. We integrate the previous method with a bootstrap algorithm Mohammad (2021). By applying the bootstrap technique, we can derive confidence intervals and improve the robustness of our cross-correlation analysis. This iterative process allows us to quantify the uncertainties associated with our measurements, ultimately enhancing the statistical significance of our findings. We consider one bin in the $\mu$-variable with $\mu_{\text{max}}=1$ Sinha (2023) and 60 bins in the separation $s$-variable. The resampling number employed in the bootstrap method  [^8] It should be emphasize that the bootstrap algorithm offers several advantages in our statistical analysis. To be more precise, it provides a more robust estimate of uncertainties by allowing replacement resampling, which better captures the underlying distribution of the data. It is particularly beneficial when the quasar pair dataset size is small compared to the larger quasar sample, ensuring that we account for variability and improve our confidence in the derived results. is set to $N_{\text{rs}}=100$  [^9] We investigated various values of the resampling number $N_{\text{rs}}$ and found that the results remain largely invariant, showing no significant variation across the different configurations tested., consistent with the approach outlined in Mohammad (2021).
@@
Fig. (13) illustrates the cross-correlation function on different scales. The analysis reveals that the cross-correlation signal is considerably stronger for separations $s<20\text{Mpc}h^{-1}$. In other words, quasar pairs within this range exhibit a more significant clustering tendency, suggesting a closer spatial relationship among these objects. The pronounced signal at this separation may imply underlying physical processes or structures influencing the distribution of quasars. The cross-correlation function of (quasar pairs)$\times$(quasars) is found to be 5.98 times higher than that of the autocorrelation function of quasars when considering separations $s<20\text{Mpc}h^{-1}$. At large separations, both signals exhibit a similar decline, demonstrating a consistent behavior in the correlation functions. To gain insight into how the correlation length for the cross-correlation of quasar pairs with quasars compares to the quasar-quasar case, we can observe that the condition $\xi_{pq}(s_{0,pq})=1$ corresponds to a scale of approximately $s_{0,pq}\simeq 18.94$. In contrast, for the autocorrelation case, where $\xi_{qq}(s_{0,qq})=1$, the scale is around $s_{0,qq}\simeq 9.64$. The visual examination indicates a significant difference in the correlation lengths between these two scenarios, highlighting the influence of the cross-correlation on the spatial distribution of quasar pairs.  [^10] To grasp the significance of our earlier results, we can consider the formulation of the signal-to-noise ratio in the context of cross-correlation: $S/N\propto\sqrt{N_{q}N_{p}}\xi_{pq}(s)$. The latter relationship emphasizes that a nonzero pair detection is fundamentally encoded in the amplitude of the cross-correlation function, $\xi_{pq}(s)$, when juxtaposed against the background of the quasars Menard & Bartelmann (2002). Here, the interaction between the number of quasars $N_{q}$ and the number of potential counterparts $N_{p}$ amplifies the sensitivity of our measurements, allowing us to discern genuine correlations amidst the noise inherent in the quasar distribution. As expected, the correlation length of the quasar pairs is greater than that of individual quasars.

### v2

@@ 4 Quasar Pair Catalog
We also examine various bounds of $|\Delta V|$ to investigate how the number of pairs detected changes as this condition is relaxed within the redshift interval $z\in[1.2,2.8]$. For the condition $|\Delta V|\leq 600\text{km}/s$, the number of pairs detected decreases to 1915, representing $39\%$ of the total. In contrast, when we consider $|\Delta V|\leq 1200\text{km}/s$, the number of pairs detected increases to 3605, which represents $73\%$ of the entire subsample. Consequently, we will focus on the most conservative case moving forward, which corresponds to $|\Delta V|\leq 900\text{km}/s$. The distribution of quasar pairs within the RA-DEC plane is illustrated in Fig. (6) for the conservative scenario.
Figure 6: For the subset of 2777 quasar pairs that meet the criterion $|\Delta V|\leq 900\text{km}/s$, a scatter plot illustrating these pairs is shown. This visualization incorporates a normalized color bar that conveys the variation of $\Delta z$ among the quasar pairs, effectively representing this continuous variable through a color scale.
We address another observational point regarding the detection of quasar pairs. Firstly, we pose a fundamental question: Is the SDSS camera capable of effectively resolving two quasars that constitute a confirmed pair? This query is essential for understanding the limitations of observational data and the ability of the SDSS to distinguish quasar pairs, which is crucial for accurate pair identification and subsequent analysis. To answer that question, we use a conservative angular resolution for the SDSS telescope Lyke et al. (2020), with $\theta\geq\theta_{SDSS}=5^{{}^{\prime\prime}}$, and a mean redshift value of $z_{\text{mean}}=2.4$. The physical distance corresponding to the SDSS angular resolution yields approximately $d^{\text{min}}_{p}\simeq 0.03\text{Mpc}$ [cf. Fig. (5)]. Such findings are significant in light of previous observations, which indicated that the distribution of distances between quasar pairs is around $d=0.986\text{Mpc}$. Thus, we can conclude that the SDSS camera is capable of resolving these objects.
Another important question is whether the quasar pair sample might be contaminated. For example, background galaxies can create multiple images of a single quasar, making it difficult to identify true quasar pairs. To tackle this issue, we can reassess the quasar pair sample by focusing only on pairs with $\Delta V\leq\sigma^{SDSS}_{V}$, where $\sigma^{SDSS}_{V}(z)=60\rm{km}/s$ represents the error in the spectroscopic measurements from the SDSS DR16 catalog Lyke et al. (2020), and $\Delta V=c(z_{2}-z_{1})$.Our analysis revealed that out of the 2777 detected quasar pairs, only $9.9\%$ could be potentially contaminated. To investigate these cases further, we used the specific coordinates of these quasars to examine their corresponding spectra from the SDSS DR16 catalog. However, it’s important to note that the SDSS DR16Q catalog has already looked into several contamination issues within the full sample, including duplicity due to the observational strategy, reporting up to $1.3\%$ of potential contamination Lyke et al. (2020). Our criteria for assessing contamination are as follows: If both quasars in a pair show identical features in their spectra, it suggests they may be different images of the same quasar, influenced by foreground galaxies. On the other hand, if the elements of a quasar pair exhibit different amplitudes and positions of their emission lines, we can confidently conclude they are not multiple images of the same object. The position of each object in the sky and its spectrum can be accessed through the skyserver-dr16 wesbsite  [^7] https://skyserver.sdss.org/dr16/en/tools/chart .
[…]
To further characterize the quasar pair catalog, we perform a numerical estimation of the comoving number density of quasars and quasar pairs within the redshift interval of $[1.2,2.8]$. The fraction of the sky surveyed is approximately $0.65\%$ covering $26732\text{deg}^{2}$, leading to a numerical (comoving) density of $n_{q}\simeq 5.8\times 10^{-8}\text{Mpc}^{-3}$. This value aligns well with the findings reported in Croom et al. (2005); namely, $n_{q}\simeq 8.2\times 10^{-9}\text{Mpc}^{-3}$. Taking into consideration a minor sky fraction for the quasar pairs $f_{\text{sky, pair}}=0.54$, we observe a notable reduction in the estimated numerical (comoving) density of these pairs. It approximately yields $n_{\text{pair}}\simeq 2.4\times 10^{-10}\text{Mpc}^{-3}$. The latter finding highlights the significant influence that observing a restricted area of the sky, combined with a relatively small sample of detected pairs, has on the estimated density of quasar pairs, particularly in contrast to the densities associated with individual quasars Croom et al. (2005).
5 Luminosities and Color-Color Maps
In this section we explore wether quasar pairs exhibit particular properties with respect to quasars in isolation. Here we focus on the luminosity of the quasar pairs Wu&Shen (2022), as well as the luminosity ratio of the pair members.
Quasar luminosties must take into account the K-correction, Ross et al. (2013). Our approach follows Lyke et al. (2020), using a pre-calibrated K-correction table from Richards et al. (2006). The K corrected absolute magnitude in the I-band is calculated using the relations:
$m_{I,\text{ap}}=M_{I,\text{abs}}+DM(z)+K(z)+A(z),$ (3)
where $m_{I,\text{ap}}$ is the apparent magnitude, $M_{I,\text{abs}}$ is the absolute magnitude in the I-band, $DM(z)=\log_{10}\left[D_{L}(z)/10\rm{pc}\right]$ is the distance modulus, and $A(z)$ represents extinction. The K-correction, found in the second-to-last term of the equation, can be specified at $z=2$ as detailed in Appendix B of Ross et al. (2013). Our method for integrating K-correction closely follows the procedures from Ross et al. (2013) and Lyke et al. (2020). We use the Planck 2018 fiducial cosmology for distance calculations. The absolute magnitude $M_{I,\text{abs}}$ for a given quasar is obtained by correcting the apparent magnitude for extinction, distance modulus, and K-correction. Apparent magnitudes in the I-band are taken from the PSFMAG column in the FITS file of the SDSS DR16 QSO catalog Lyke et al. (2020) [^8] https://www.sdss4.org/dr16/algorithms/qso_catalog. K-corrections come from Richards’ table, which is limited to a specific redshift range Richards et al. (2006). For each redshift $z$, we find the closest entry in the K-correction table to extract the corresponding K-correction value. This process allows us to calculate the corrected absolute magnitude $M_{I,\text{abs}}$ for each quasar. Once we have $M_{I,\text{abs}}$, we can convert it to luminosity in the I-band in solar units using:
$L=10^{-0.4(M_{I,\text{abs}}-M_{I,\odot})}L_{\odot},$ (4)
@@ 5 Luminosities and Color-Color Maps
where $L_{\odot}$ is the solar luminosity and $M_{I,\odot}$ is the absolute magnitude of the Sun in the I-band. Thus, we obtain reliable estimates of the luminosities in the I-band allowing for further analysis of quasar in pairs and in isolation. In Fig. (7), we display the distribution of luminosities, showing that $L=(1.75\pm 0.52)\times 10^{16}L_{\odot}$.
Figure 7: Histogram of the quasar luminosity distribution after applying the K-corrections. The horizontal axis is displayed on a logarithmic scale.
It is now well understood that galaxy evolution, shaped by AGN feedback, often leads to quasars residing in low-mass dark matter halos Fanidakis et al. (2013). Furthermore, the number density of quasars per unit luminosity follows a double power-law relationship within the redshift range $z\in[0.3,2.2]$, while showing log-linear behavior for $z\geq 2.2$ Ross et al. (2013). This suggests that quasar pairs-acting as proxies for protoclusters Muldrew et al. (2015), Alberts &\&& Noble (2022)-may share similarities due to their common dark matter halo.
Here we aim at addressing if quasar pairs within the same dark matter halo represent a distinct class of objects, rather than just a simple combination of two individual quasars from the sample. To address the latter point, we examine the distribution of luminosity ratios, $L_{1}/L_{2}$, for the quasar pairs, where $L_{1}$ is the smaller luminosity and $L_{2}$ is the larger one. This analysis will allow us to investigate the intrinsic relationships between the luminosities of these unique objects, offering insights into their relative behaviors. Additionally, we will generate 100 random sets of control pairs derived from the isolated quasar set. These ensembles will not satisfy the typical constraints of $\Delta V\leq 900\text{km}/s$ and $d_{p}\leq 2\text{Mpc}$  [^9] A similar analysis was carried out on 500 random sets of synthetic pairs generated from the isolated set, showing no significant deviations.. This approach will provide a baseline for comparison, which is essential for understanding the characteristics of the quasar pairs and their luminosity distributions in relation to the control pairs. By contrasting the observational data with these control constructs, we aim to identify any significant patterns or anomalies that may arise, thereby enriching our comprehension of these astrophysical entities.
Figure 8: Distribution of the luminosity ratios for true pairs and control ones. The true quasar pairs are indicated in blue, whereas the control pairs are shown in gray.
We construct the control pair sample imposing that the angular separation between the two objects exceed $5.7^{\circ}$ (1 radian), and the maximum allowable redshift difference is restricted to $\Delta z\leq 0.003$. We also verified that the histograms of $\Delta z$ for the true pairs and control pairs are quite similar. This consistency in their distribution assures that there is no significant difference in the redshift differences between the two groups, indicating that the selection criteria for both true and control pairs are comparable. We generated 500 sets of control pairs, each containing the same number of detected pairs, resulting in 615 control pairs per set. We focused on assembling the new set using a random algorithm to avoid any selection bias  [^10] While our focus is on the redshift interval $z\in[2.2,2.8]$, it is noteworthy that a similar analysis has been conducted across other redshift bins, yielding comparable results..
Fig. (8) illustrates the differences between the true quasar pairs and the control pairs. The mean luminosity ratio $L_{1}/L_{2}$ shows significant variations between these two datasets. When we calculate the fraction of real pairs that meet the condition $F_{\text{true pair}}(L_{1}/L_{2}>0.75)$ and compare it to the fraction for the control pairs $F_{\text{control pair}}(L_{1}/L_{2}>0.75)$, we can determine the ratio of these fractions, defined as ${\cal{R}}=F_{\text{true pair}}/F_{\text{control pair}}$. The goal of calculating $F_{\text{control pair}}(L_{1}/L_{2}>0.75)$ is to see if the control pairs can mimic the histogram of quasar pairs that show $L_{1}/L_{2}>0.75$.
The results show that ${\cal{R}}>1$, highlighting that the two sets of pairs have fundamentally different properties, likely due to the different accretion mechanisms affecting the luminosity ratios $L_{1}/L_{2}$. Our analysis finds that $62.32\%$ of true pairs exhibit $L_{1}/L_{2}>0.75$, while only $57.56\%$ of control pairs meet this criterion. Thus, the calculated value of ${\cal{R}}=1.08$ indicates that true pairs show a luminosity ratio $L_{1}/L_{2}>0.75$ compared to their control sample counterpart. Moreover, our findings are consistent across all 500 datasets of control samples, each containing 615 pairs, reinforcing the robustness of our conclusions.
@@
Figure 9: Distribution of the logarithm of the sum of luminosities for true pairs and control pairs. The true quasar pairs are indicated in blue, whereas the control pairs are shown in gray.
Another approach to investigate the validity of random-generated data sets compared to true pairs is to analyze the behavior of the total luminosity of the pair, i.e., the sum of the member luminosities. This is motivated by the possibility of a different behavior for quasars in association with respect to randomly selected objects (notice that the results of the luminosity ratio analysis could imply both quasars to have either low or high luminosities). As it can be seen in Fig. (9) the resulting distribution of the logarithm of the sum of the luminosities differs for true pairs and control pairs. The histogram of the true pairs shows a narrower spread around its mean, indicating that these pairs are more closely clustered in terms of their luminosity ratios. In contrast, the histogram for the control pairs exhibits a broader spread, indicating a larger luminosity variability among them. This difference in distribution reinforces the notion that true pairs are more consistent in their luminosity characteristics in contrast to the control sample pairs tend which display a wider range of luminosity ratios.
Our analysis shows a significant difference between the two datasets, suggesting that the mechanisms driving the luminosity characteristics of quasars in association are intrinsically distinct. These differences may largely arise from the co-evolution of quasars with their surrounding environments, which influences their accretion processes and, in turn, their luminosity profiles Porciani et al. (2004). Such interactions with the environment can significantly shape the properties of quasars, leading to observable variations in luminosity distributions Eftekharzadeh et al. (2015).
@@
Figure 10: The distribution of luminosity ratios for true pairs and simulated pairs is illustrated. The true quasar pairs are represented in blue, while the simulated pairs are depicted in gray.
We aim to further quantify the differences observed between true quasar pairs and their control counterparts. The main goal is to evaluate whether a modified set of pairs from the control sample, hereafter named ”simulated” quasar pairs, can effectively approximate the real data. In this approach, we focus on the tercile of the control sample associated with the less luminous quasar and increase its luminosity by $30\%$.
Figure (10) shows that this procedure creates a simulated quasar pair sample with a distribution of luminosity ratios $L_{1}/L_{2}$ that closely resembles the actual data, showing the strength of similarity effect. Fundamentally, this observation indicates that true pairs exhibit closely similar luminosity values. In contrast, control pairs show a more pronounced disparity between their most and least luminous members. This distinction may imply that reinforcing the hypothesis of a common cosmic evolution of our pair sample.
@@
Figure 11: Top panel: Color-color map for the entire subsample of quasars. Bottom panel: Color-color map for the detected quasar pairs. Here, $\log(N)$ represents the logarithm of the number of quasar pairs that fall within each region of the color-color map. Notice that the maximum or minimum values may be outside the range of displayed data points in some cases.
@@
In the final stage of characterizing the quasar pairs, we perform a comparative analysis of the color-color map of these pairs against that of the entire selected sample. As illustrated in Fig. (11), the color-color map derived from the PSF magnitudes in the $g$, $i$, $r$ and $z$ bands for the complete subsample reveals a range of values given by $-6.04\leq(g-r)\leq 5.64$ and $-3.68\leq(r-z)\leq 7.15\leavevmode\nobreak$ Lyke et al. (2020). However, when we focus on the detected pairs, these maximum and minimum values are significantly constrained, $-1.55\leq(g-r)\leq 2.21$ and $-1.17\leq(r-z)\leq 2.53$. It is important to note that $90\%$ of the data points associated with the quasar pairs fall within the interval $(r-z)\times(g-r)\in[-0.10,0.85]\times[-0.17,0.42]$. The latter results indicate older stellar populations or dust effects along with the possibility of bluer colors signifying ongoing star formation. This is consistent with the current understanding of quasar environments and their host galaxies Kauffmann &\&& Heckman (2009). For the $90\%$ of the detected quasar pairs, we obtain $(i-z)\times(r-i)\in[-0.06,0.69]\times[-0.13,0.30]$. Once again, the negative minimum values suggest the presence of younger, hotter stars, while the positive maximum values indicate the influence of older stars or dust obscuration. These interpretations are well supported within the framework of quasar studies Richards et al. (2006). Different sequences of color-color maps are displayed in Fig. (12)
@@ 6 Cross-correlation Quasar-Pair
To ensure that the identified pairs are indeed physical composite objects located within the same dark matter halo and are likely experiencing similar accretion processes—rather than being merely coincidental nearby quasars—we will implement a series of verification tests. Among these, we will conduct an analysis of the cross-correlation of signals between the quasar sample and the corresponding subsample of pairs, as we stated before. We implement the same protocol as previously done. Using the corrfunc package Sinha & Garrison. (2020), we calculate $\xi(s,\mu)$, subsequently converting this to the number of pairs with the assistance of the convert-3d-counts-to-cf routine. That enables us to derive the Landy & Szalay (1993) estimator for the cross-correlation,
@@
where $D_{q}$ represents the set of quasars belonging to the subsample, and $D_{p}$ denotes the dataset of the quasar pairs, using the mid-vector between the two quasars to represent each pair. We implement a method to generate an associated random set that satisfies the specifications outlined in the corrfunc manual Sinha (2023). Specifically, we ensure that the size of the random set and the data set is consistent with the relationship $N(R_{p})=3N(D_{p})$. The latter approach improves statistical robustness and mitigates any potential biases in our correlation measurements. We integrate the previous method with a bootstrap algorithm Mohammad & Percival (2021). By applying the bootstrap technique, we can derive confidence intervals and improve the robustness of our cross-correlation analysis. This iterative process allows us to quantify the uncertainties associated with our measurements, ultimately enhancing the statistical significance of our findings. We consider one bin in the $\mu$-variable with $\mu_{\text{max}}=1$ Sinha (2023) and 60 bins in the separation $s$-variable. The resampling number employed in the bootstrap method  [^11] It should be emphasize that the bootstrap algorithm offers several advantages in our statistical analysis. To be more precise, it provides a more robust estimate of uncertainties by allowing replacement resampling, which better captures the underlying distribution of the data. It is particularly beneficial when the quasar pair dataset size is small compared to the larger quasar sample, ensuring that we account for variability and improve our confidence in the derived results. is set to $N_{\text{rs}}=100$  [^12] We investigated various values of the resampling number $N_{\text{rs}}$ and found that the results remain largely invariant, showing no significant variation across the different configurations tested., consistent with the approach outlined in Mohammad & Percival (2021).
To begin, we aim to illustrate the overall behavior of the cross-correlation amplitude across the entire redshift range of interest, $1.2\leq z\leq 2.8$. Subsequently, we will revisit this cross-correlation function under various scenarios, which will involve adjusting the redshift bins and exploring different combinations of cuts based on $L_{1}/L_{2}$ and $\log_{10}\big{(}L_{1}+L_{2}\big{)}/L_{\odot}$. The aforementioned cuts are well-founded, as they are supported by the luminosity analysis of quasar pairs presented in section 5.
In Figure (13), we present the cross-correlation function for this specified interval, comparing its signal to both the autocorrelation of all quasars and that of isolated quasars—those not associated with any pairs. Our analysis indicates that the cross-correlation between the pairs and isolated quasars does not produce a stronger signal than the overall cross-correlation derived from the complete sample within this interval. We observe that the cross-correlation signal between pairs and quasars consistently surpasses the autocorrelation signal of quasars by a significant margin. The signal is clearly stronger than the auto-correlation for separations of $s<25\text{Mpc}\text{h}^{-1}$. While both correlation signals exhibit a similar decline at larger separations, as illustrated in the log-log plot in Fig. (13), the correlation scale for quasar pair correlations is approximately $s_{0,pq}\simeq 21.41$, nearly doubling the correlation length of the quasar sample, which is $s_{0,qq}\simeq 12.25$. The latter observation indicates that quasar pairs within this range demonstrate significant clustering, implying a stronger spatial association influenced by underlying physical processes. Such clustering is one of the reasons why quasar pairs can be regarded as proxies for protoclusters Muldrew et al. (2015), Alberts &\&& Noble (2022).
@@
Now, we focus on the cross-correlation function on different scales, $2.2\leq z\leq 2.8$ [cf. Fig. (14)]. The analysis reveals that the cross-correlation signal is considerably stronger for separations $s<20\text{Mpc}h^{-1}$. In other words, quasar pairs within this range exhibit a more significant clustering tendency, suggesting a closer spatial relationship among these objects.. The cross-correlation function of (quasar pairs)$\times$(quasars) is found to be 5.98 times higher than that of the autocorrelation function of quasars when considering separations $s<20\text{Mpc}h^{-1}$. At large separations, both signals exhibit a similar decline, demonstrating a consistent behavior in the correlation functions. To gain insight into how the correlation length for the cross-correlation of quasar pairs with quasars compares to the quasar-quasar case, we can observe that the condition $\xi_{pq}(s_{0,pq})=1$ corresponds to a scale of approximately $s_{0,pq}\simeq 18.94$. In contrast, for the autocorrelation case, where $\xi_{qq}(s_{0,qq})=1$, the scale is around $s_{0,qq}\simeq 9.64$. The visual examination indicates a significant difference in the correlation lengths between these two scenarios, highlighting the influence of the cross-correlation on the spatial distribution of quasar pairs.  [^13] To grasp the significance of our earlier results, we can consider the formulation of the signal-to-noise ratio in the context of cross-correlation: $S/N\propto\sqrt{N_{q}N_{p}}\xi_{pq}(s)$. The latter relationship emphasizes that a nonzero pair detection is fundamentally encoded in the amplitude of the cross-correlation function, $\xi_{pq}(s)$, when juxtaposed against the background of the quasars Menard & Bartelmann (2002). Here, the interaction between the number of quasars $N_{q}$ and the number of potential counterparts $N_{p}$ amplifies the sensitivity of our measurements, allowing us to discern genuine correlations amidst the noise inherent in the quasar distribution. As expected, the correlation length of the quasar pairs is greater than that of individual quasars.